OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. cos x - sin squaredx - 1

OpenStudy (anonymous):

Remember that sin2x = 2sinxcosx so cosx - 2sinxcosx = 1. factor the cosx

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

o no.. was the expression \[cosx - \sin ^{2} x -1\]

OpenStudy (anonymous):

In that case \[\sin ^{2}x = 1 - \cos ^{2}x\]